OpenStudy (anonymous):

1-1/5X I=3 *sn there is a line in front of the 1 and behind the 5X

OpenStudy (amistre64):

|-1/5x| = 3 ? like that

OpenStudy (anonymous):

yes :)

OpenStudy (amistre64):

x = 15 or x = -15 15/5 = 3

OpenStudy (anonymous):

okay so its 15 and -15

OpenStudy (amistre64):

yes, since its an absolute value; both are solutions

OpenStudy (anonymous):

can i ask another one?

OpenStudy (amistre64):

sure

OpenStudy (anonymous):

solve the inequality I 4-xI <8 the is a line under the sigh

OpenStudy (anonymous):

this sigh <

OpenStudy (amistre64):

<= is what that line means with it

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

4-x <= 8 and -(4-x) <=8 -x <= 4 x <= 12 x >= -4 and x <= 12 your solution interval is [-4,12]

OpenStudy (anonymous):

oh okay thanks so much

OpenStudy (amistre64):

youre welcome :)